OpenStudy (anonymous):

What is % 15? So if I have 129, or 2501, how do you get the answer? 129 % 15 etc?

OpenStudy (anonymous):

If you want to find out what a certain percentage of a number is, you change the percentage number into a decimal. For example, 15% would be .15 and 16% would be .16, and so on. Then, you multiply this decimal by the number you want to find out what the percentage of the number is. So you would do .15 x 129 = 19.35, or .15 x 2501 = 375.15.

OpenStudy (anonymous):

this website could help you for further info http://www.themathpage.com/arith/what-percent.htm

OpenStudy (anonymous):

Ah I was looking at the question and the answers were provided. I didn't know what they meant. % is mod. But I get it now, for example: 129 mod 15 = 15 * 8 = 120 129 - 120 = answer is 9.

OpenStudy (anonymous):

mod is the remainder you get when you divide the first number by the second number.

OpenStudy (anonymous):

It looks like you just figured that out though :) Well done!